=At 1650°C, the equilibrium constant for the reaction CO(g) + 2H 2 (g) =CH 3 OH(g) has the value K p = 1.4 x 10 –7 . Calculate G° for this reaction at 1650°C.
Answer #1

Ans is

304.1 kJ/mol

for co t 2 H₂ CHOH (9) (9) (9) (1+2) on - = -2 kp ke - (R7) on where R = = 8.314 Tank 1650°C T- 1923k - li4 X10 xiol ke - -2-16 Kc 5,5x10 Now Aga RT enk whore In K = dn (s. 5x106) -19, 0 sĞ = 68.314)x(923) * (-19.0) - 304131.431 my = 304. KJ mๆ KJ A